OpenStudy (mven):

Find the radius of a circle whose area equals the area of a rectangle that measures 2 ft. by 11 ft. (Use pi = 22/7)

OpenStudy (anonymous):

Area of rectangle = 2 * 11 = 22 square feet Area of circle = pi * r^2 pi * r^2 = 22 r^2 = (22 * 7)/22 = 7 r=7√

OpenStudy (agl202):

https://answers.yahoo.com/question/index?qid=20100519174450AA131nn @your_mom017 LOL

OpenStudy (mven):

lol popular question i guess. thanks

OpenStudy (anonymous):

what is that

OpenStudy (agl202):

Where ur answer is from!!! >_> lol

OpenStudy (mven):

how about this one? A circle has a radius of 8 inches. Find the area of a sector of the circle if the sector has an arc that measures 45°

OpenStudy (anonymous):

i didnt get my answer from there im actually in math class right now

OpenStudy (mven):

whats with the??????????

OpenStudy (agl202):

I Know right! lol @your_mom017

OpenStudy (anonymous):

You should start by converting the angle 45 into radian.. then formula for area is = 0.5*r^2* theta.. theta is the angle here.. you have the radius too given in the question.. so just plug in the values. The formula for finding the length of the sector: s=rtheta For the area of a sector: A=(1/2)r2θ

OpenStudy (mven):

would that be 8pi?

OpenStudy (anonymous):

should be

OpenStudy (mven):

awesome thanks. one more Find the approximate area of a circle with radius equal to 8 ft. (Use = 3.14)

OpenStudy (mven):

200.96?

OpenStudy (anonymous):